An economy has 200 million people divided among the following groups; 50 million have full time jobs, 20 million have one part-time job, 40 million have two part-time jobs, 30 million would like to work and are looking for jobs, 10 million would like to work but are so discouraged they have given up looking, 10 million are running their own business, 10 million are retired, and 30 million are small children.

a. What is the size of the labor force?

b. What is the labor force participation rate?

c. What is the unemployment rate?

d. Why might the unemployment rate be a poor indicator of how well a country's labor resource is being utilized?
Answer #1

Answer:

a. Size of the labour force:

50 million have full time jobs + 20 million have one part-time job + , 40 million have two part-time jobs + 30 million would like to work and are looking for jobs + 10 million are running their own business (50+20+40+30+10) = 150 million

b. labor force participation rate:

LFPR = Labor Force / Civilian Non-Institutionalized Population

Labor force=150 million

Civilian Non-Institutionalized Population= Labor Force+ 10 million are retired+10 million would like to work but are so discouraged they have given up looking = 170 million

LFPR = 150/170 *100= 88.24%

c. unemployment rate: 88.24 %

labor force participation rate and unemployment rare both are same.

d. Why might the unemployment rate be a poor indicator of how well a country's labor resource is being utilized?

The unemployment rate is a poor indicator of how well a country’s labor resource is being utilized because

  • it doesn’t include discouraged Work forces
  • It Overlooks Other Marginally Involved Workers
  • No proper separation of Part-Time and Full-Time Workers
  • It Doesn’t Consider Whether People Have Low-Paying Jobs or high paying jobs
  • It Doesn’t reflect the Long-Term Unemployment Rate